MHATT-CAT Useful Operational Data for Run 3 of FY02
Eric Dufresne, Dohn Arms, Don Walko
MHATT-CAT
Started July 19, 2002
(www.mhatt.aps.anl.gov/Sectors/Sector7/Operations/FY02_run3/)
Table of Content:
Introduction.
The work shown below is an account of useful operational data taken during
run 3 of FY02 on 7ID. This page will contain useful stability information on the
7ID High Heat Load Monochromator and also on the various repairs and problems
identified during the run.
Thursday, July 18, 2002. Time series with a 0.4 mm (H) by 0.4 mm (V)
white beam, with the mono at 10.0 keV.
The mono is stabilizing its second crystal temperature after the gap was opened
from around 11 mm to 21.768 mm around 16h12 on 7/18/02. The data was started
on 17h28 and lasts for more than 16 hours. Three figures are shown below.
Fig. 07-18.1 shows a time series of the intensity not
normalized to beam current decay. the data was taken every second. Matt DeCamp
tweaked the second crystal picomotor several time during the 8 hours. I am now
pretty convinced that the picomotor mechanical stage pth drifts after a tweak
I think in the future we should use the piezo electric transducer to tweak the
second crystal. It would make the second crystal more stable.
Fig. 07-18.2 shows a time serie of the 7ID-C Beam
position. The vertical beam motion is very large during the first half of the
time series (ignore large drops due to hutch access). Note also that no data is
recorded when someone enters in 7ID-C.
Fig. 07-18.3 shows a time serie of the second crystal
temperature during the same time serie. The temperature takes about 4-5 hours
to stabilize after opening the gap. I don't believe the data after 10 hours.
I suspect noise pick-up which I vaguely recall seems to be worse at night.
Fig 07-18.1. Time series of the beam intensity in 7ID-C, starting at
17h28 on 07/18 and lasting 16 hours. The 7ID-C diode sum, the ring current,
ion chamber in B and C are shown.
Fig 07-18.2. The beam position, 49 m from the source, or 19 m from the
High Heat Load mono in 7ID-C during the same time series as
Fig. 07-18.1. The time series is started on 07/18 at 17h28, and lasts
for 16 hours. The beam is not stable in the first half of the time series due to
thermal drift of the monochromator and possible mechanical drifts from the
second crystal (pth).
Fig 07-18.3. The second crystal temperature during the same time series as
Fig. 07-18.1. The time series is started on 07/18 at 17h28, and lasts
for 16 hours.
Tuesday, August 14, 2002. Time series with a 0.5 mm (H) by 0.5 mm (V)
white beam, with the mono at 10.0 keV.
This time serie was started on 8/14 at 01h07 and lasts for 8.5 hours. The L5-20
opening was set to 0.5x0.5 mm shortly after the start of the time serie, within
the first three minutes. They were set at 0.3x0.3 mm before so the white beam
area of the first crystal increased by (0.5/0.3)^2 ~ 2.8, a significant power
increase.
Fig. 08-14.1 shows a time serie of the intensities
in 7ID-B, and C as well as the diode sum. The intensity drifts for several
hours after the change of L5-20 opening. Fig. 08-14.2
shows a time serie of the beam position in the same time serie. The horizontal
position rapidly change by 20 um in about one hour, but the vertical position
drifts by more than 110 um over 5 hours, following the trends of Fig. 08-14.1.
The second crystal pth was not tweaked, so no mechanical instabilities can
explain the vertical drift seen. The beam motion and loss of intensity is
likely due to the change in temperature of the second crystal.
Fig 08-14.1. Time series of the beam intensity in 7ID-C, starting at
01h07 on 08/14 and lasting 8.5 hours. The 7ID-C diode sum, the ring current,
ion chamber in B and C are shown.
Shortly after the start of the time
series, the L5-20 slit opening was increased to
0.5x0.5 mm from 0.3x0.3mm.
Fig 08-14.2. The beam position, 49 m from the source, or 19 m from the
High Heat Load mono in 7ID-C during the same time series as
Fig. 08-14.1. The time series is started on 08/14 at 01h07, and lasts
for 8.5 hours. The beam drifts in the vertical direction because the second
crystal temperature is changing after opening the L5-20 opening.
The second crystal picomotors
were not tweaked after the opening of the slits.
Last modified on 10/22/2002
by Eric Dufresne,
Bldg 432 Rm D007 630-252-0274
|